OpenStudy (anonymous):

If the volume of a box is 1,300 cubic inches and all its dimensions are equal, what is the side length of the box in inches? The formula is (V = lwh).

hartnn (hartnn):

all sides equal, so V=l^3 so l = V^(1/3) or cube root of V

hartnn (hartnn):

so, cube root of 1300 is ?

OpenStudy (anonymous):

lol...sorry got distracted by chat... umm the square root of 1300=about 11..correct? 10.9

hartnn (hartnn):

absolutely! 10.9 inches, to be complete :)

OpenStudy (anonymous):

yayy! thank you! :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

lol that was so simple and easy! :)

hartnn (hartnn):

yes, it was !

OpenStudy (anonymous):

sorry...

hartnn (hartnn):

for what ?

OpenStudy (anonymous):

sking a stupid an easy question..

hartnn (hartnn):

no problem at all :)

OpenStudy (anonymous):

ok, thank you! :)